Tutor,
I am sorry this is not from the book, I just need to know if I wrote this correctly please help.
Translate the following statement into an algebraic equation. Let x represent the number. 3 less than 20 times a number is 9 times that same number.
I came up with 3-20x=9x
Is this the proper equation for this sentence? Thank you inadvance for your help.

Found 2 solutions by jim_thompson5910, checkley75:
Answer by jim_thompson5910(28715)

When you take 3 less of something, you are subtracting 3 from it. So if you take 3 away from 20 times a number it should look like this:

20x-3=9x

Answer by checkley75(3666)

20X-3=9X
20X-9X=3
11X=3
X=3/11 ANSWER.
PROOF
20(3/11)-3=9(3/11)
60/11-3=27/11
60/11-33/11=27/11
(60-33)/11=27/11
27/11=27/11